OpenStudy (happykiddo):

DIFF EQ: Is y = sin t a solution to (dy/dt)^2= 1 − y^2? Justify.

OpenStudy (prettygirl_shynice):

ill help

OpenStudy (solomonzelman):

Find the second derivative of sin(t).

OpenStudy (solomonzelman):

The reason I am asking you this, so that you can plug in the second derivative of sin(t) instead of (dy/dt)^2, and to plug in sin(t) instead of y.

OpenStudy (prettygirl_shynice):

oh

OpenStudy (prettygirl_shynice):

okay hold on @jigglypuff314

OpenStudy (prettygirl_shynice):

im srry

OpenStudy (solomonzelman):

Just to clarify in case you didn't know (dy/dt)^2 is denoting the second derivative of the function y, with respect to t.

OpenStudy (solomonzelman):

sorry, but what does omm stand for? (not very good with English slang)

OpenStudy (prettygirl_shynice):

oh it mean didntt mean to say that

OpenStudy (solomonzelman):

For example I have the following differential equation, \(\color{#000000 }{ \displaystyle y''+2y'+y=2x+4 }\) And I want to verify whether or not \(y=2x\) is a solution to the differential equation, than this is what I would do. I know that, \(\color{#000000 }{ \displaystyle y=2x }\) \(\color{#000000 }{ \displaystyle y'=2 }\) \(\color{#000000 }{ \displaystyle y''=0 }\) And I will plug this into my differential equation, \(\color{#000000 }{ \displaystyle (0)+2(2)+(2x)=2x+4 }\) \(\color{#000000 }{ \displaystyle 4+2x=2x+4 }\) So therefore, i have verified that \(\color{#000000 }{ \displaystyle y=2x }\) is a solution to \(\color{#000000 }{ \displaystyle y''+2y'+y=2x+4 }\).

OpenStudy (happykiddo):

Thanks guys you were a big help when tried to solve for the equation my solution would always end up being y=-sin(t). I see what I did wrong now.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y''=1-y^2 }\) If your solution is \(\color{#000000 }{ \displaystyle y=a\sin(t) }\) then, \(\color{#000000 }{ \displaystyle y=a\sin(t) }\) \(\color{#000000 }{ \displaystyle y'=a\cos(t) }\) \(\color{#000000 }{ \displaystyle y''=-a\sin(t) }\) and we plug that in, \(\color{#000000 }{ \displaystyle y''=1-y^2\quad \Longrightarrow \quad -a\sin(t) =1-(a\sin(t))^2 }\) and when you simplify you get, \(\color{#000000 }{ \displaystyle -a\sin(t) =1-a^2\sin^2(t) }\) \(\color{#000000 }{ \displaystyle a^2\sin^2(t) -a\sin(t)-1 =0}\) and that is clearly not the solution

OpenStudy (happykiddo):

I know the solution is y = -sin(t), but how can I show this using the first order separable ODE method. Using the N(y)dy = M(x)dx , I know that N(y) = 1/(sqrt(-y^2+1)), but why?

OpenStudy (happykiddo):

Where'd the square root come from?

OpenStudy (happykiddo):

I know that I could show that y= -sin(t) by using the method you described but how could I do it with the First order separable method? Big thanks

OpenStudy (solomonzelman):

If you differential equation, actually, is \(\color{#000000 }{ \displaystyle (y')^2=1-y^2 }\) (although I haven't ever seen derivatives raised to a power in differential equations) Then, \(\color{#000000 }{ \displaystyle y'=\pm\sqrt{1-y^2} }\) \(\color{#000000 }{ \displaystyle \pm\frac{1}{\sqrt{1-y^2}}\frac{dy}{dt}=1 }\) \(\color{#000000 }{ \displaystyle \pm\color{red}{\int}\frac{1}{\sqrt{1-y^2}}\frac{dy}{dt}\color{red}{dt}=\color{red}{\int}\color{red}{dt} }\) \(\color{#000000 }{ \displaystyle \pm\int\frac{1}{\sqrt{1-y^2}}dy=\int~dt }\) \(\color{#000000 }{ \displaystyle y=\sin \theta }\) \(\color{#000000 }{ \displaystyle dy=\sin \theta~ d\theta }\) \(\color{#000000 }{ \displaystyle\pm \int\frac{1}{\cos \theta }dy=t+c }\) and even in that case \(y(t)=\sin t\) is not a solution!

OpenStudy (solomonzelman):

I have shown why \(\rm y(t)=a\sin(t) \) is not a solution for all "a"......